OpenStudy (anonymous):

if the ratio of boys to girls in a class is 3:8 how many girls are there when there are 24 boys?

OpenStudy (bahrom7893):

24/x = 3/8 x = (24*8)/3 x = 64

OpenStudy (bahrom7893):

Please click on become a fan if I helped, I really want to get to the next level!! Thanks =)